A software company’s labor requirements currently cost $315,000/year. The labor-hour requirements are expected to increase by 12% per year over the next 5 years. If inflation is 5.0%, determine the labor costs after 5 years using:

Part a

Then-current dollars: 708511.9

Part b

Constant-worth dollars.$_______________

answer b please
Answer #1

Here,

Current Cost = $315,000

Increase per year = 12%

Inflation per year =5%

So constant worth dollars value = 315000*(1+.12)^(5) = $555137.6

Current Dollar = $555137.6 * (1+.05)^(5) =$708511.9